OpenStudy (anonymous):

find the derivative of f(x)=4x^3-1/x^2-3(3)sqrt(x)

OpenStudy (anonymous):

do you need to use , [f(x+h) - F(x) ]\h or power rule

OpenStudy (anonymous):

the power rule

OpenStudy (anonymous):

fx^3-3fx-4x^5+12x^3+1=0...idk if it's right

OpenStudy (anonymous):

i'll check

OpenStudy (anonymous):

f(x)=4x^3-1/x^2-3(3)sqrt(x) re writing question y=4x^3-1/x^2-3.x^1/3 y=4x^3-x^-2-3.x^1/3 using power rule, y'=12x^2-(-2)x^-1-3.1/3.x^-2/3 y'=12x^2+2/x-1/x^2/3